# Anole: Adapting Diverse Compressed Models for Cross-scene Prediction on Mobile Devices

### v1

@@ VI-F Real-world Experiments
As depicted in Figure 9(a), we implement all methods on the Nvidia Jetson TX2 NX connected with a 1080p HD camera to conduct real-world experiments in Shanghai city. Well-trained compressed models and the decision model are downloaded to the Jetson device. We conduct real-world experiments in seven driving scenarios with different road conditions and different time in a day. LabelImg [25] is used to label all recorded frames as the ground truth for offline analysis. Figure 10 plots the F1 score of all methods. Anole outperforms all other candidate methods in all test scenarios. We visualize the car detection results of Anole (white solid frames) and SDM (red dashed frames) in a typical night driving scenario in Figure 9(b). The inference results obtained using SDM frequently contain errors, especially false negative errors as shown in the enlarged subgraph.
VI-G Inference Latency
TABLE IV: Inference latency and memory consumption on mobile devices.
@@ VI-G Inference Latency
Nano TX2 NX Laptop Loading model Execution
$\mathcal{M}_{scene}$ + $\mathcal{M}_{decision}$ 23.2 3.1 20.8 44 584
YOLOv3 313.8 42.9 62.2 240×$\times n$1 1,730
YOLOv3-tiny 37.8 10.8 32.2 40×$\times n$ 1,120
@@
We evaluate the inference latency of the decision model $\mathcal{M}_{decision}$ and compressed models on different mobile devices. The results are shown in Table IV. The results reveal that YOLOv3-tiny exhibits significantly lower latency when compared to deep YOLOv3, which is generally deemed unsuitable for deployment on devices. For instance, the latency of YOLOv3-tiny on Jetson Nano is 87.9% lower than that of YOLOv3. This highlights the substantial potential for accelerating inference using shallow models. It is also evident that $\mathcal{M}_{decision}$ can be executed in real-time on embedded mobile devices such as Jetson Nano, with a latency as low as 23.2 ms, making it suitable for online inference applications.
VI-H Memory and Power Consumptions
@@
We investigate the memory consumption of different models from the following two aspects, i.e., loading model only, and the memory consumption during inference with a batch size of 1. Table IV demonstrates that memory consumption for loading model is significantly lower than that during inference, owing to the presence of hidden parameters during inference. We also examine the impact of different power configurations adopted by Jetson TX2 NX to the performance of Anole. The power consumption and inference speed of Anole and baselines under different power modes are shown in Figure 11, respectively. Anole achieves a 45.1% reduction in power consumption compared with SDM and a inference speed of over 30 FPS with an input power of 20W running 6 cores.
VII Related Work
VII-A DNN Prediction on Mobile Devices
To perform DNN inference on mobile devices, new DNNs are specially designed [5, 6, 7, 8] or existing DNNs are compressed to match the computing capability of a mobile device [9, 10]. First, model structure can be optimized to reduce complexity [30, 5, 6]. Second, quantization precision can be reduced to minimize computational cost, e.g., use integers instead of floating-point numbers [31, 32]. Third, the neural network model can also be accelerated by pruning, i.e., deleting some neurons in the neural network [9, 10, 33]. Scene information is also utilized for model compression on edge/mobile devices [34, 35, 36, 37]. Finally, model distillation can distill the knowledge of large models into small models [7, 8]. Such schemes ensure real-time model inference at the expense of compromised accuracy.
@@ VII-A DNN Prediction on Mobile Devices
Another direction is to divide DNNs and perform collaborative inference on both edge devices and the cloud [11, 12, 38, 39, 40, 41], or to transmit compressed sensory data to the cloud for data recovery and model inference [14, 15]. Neurosurgeon [12] partitions the computation of each DNN inference task in a layer granularity. CLIO [42] addresses the instability of network conditions and optimizes inference under different network states. These approaches need coordination with the cloud for each inference, leading to unpredictable inference delays when the communication link is unstable or disconnected. However, they prove inadequate for cross-scene mobile inference scenarios where even deep models are unable to cope.
VII-B Cross-scene DNN Prediction
@@
Data-driven machine learning models face challenges in maintaining robust inference performance when dealing with cross-scene inference [43]. One natural approach for scene partitioning is to partition the scene based on prior knowledge or historical samples. First, based on prior knowledge, a similarity graph is constructed to cluster similar domains together [44]. However, obtaining such prior knowledge based on domain expertise can be challenging. Second, the original data or their extracted features can be utilized for more automated scene partitioning [45]. However, these methods may result in the loss of critical information in complex systems [29]. Cross-scene DNN prediction can also be enhanced given a golden model in the cloud for online sample labelling [46, 47]. However, the existence of a perfect or golden model is not always feasible.
VII-C Mixture of Experts
@@
In recent years, we have witnessed the success of Mixture of Experts (MoE) [48, 49], especially in efficient training of large language models (LLM). MoE employs multiple experts for model training, each for one domain. Then, a gate network will be used to determine the correspondence between samples and experts. Though inspired by MoE, Anole differs from MoE in the following 2 aspects. First, experts in MoE are diversified by constraints of losses, but they themselves cannot be related to the scene. In fact, the main purpose of MoE is to expand the number of model parameters, rather than to customize and select scene-specific models. Second, MoE is just a model architecture, and models based on MoE architecture still need to deploy the entire model during deployment. Therefore, MoE-based models often require a significant amount of memory, which is unacceptable for mobile agents like UAVs. In contrast, Anole employs multiple compressed models for online model inference, each designed for one scene. Only a few compressed models are needed to be deployed during online inference. Therefore, Anole is more suitable for mobile devices only with limited resources.
VIII Conclusion
@@
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China (Grant No. 61972081), the Natural Science Foundation of Shanghai (Grant No.22ZR1400200), and the Fundamental Research Funds for the Central Universities (No. 2232023Y-01).

### v2

@@ VII-F Real-world Experiments
TABLE VI: Inference latency and memory consumption on mobile devices.
@@
Nano TX2 Lap. Loading Execution
$\mathcal{M}_{sce.}$ + $\mathcal{M}_{dec.}$ 23.2 3.1 20.8 44 584
[…]
@@
VII-G Inference Latency
We evaluate the inference latency of the decision model $\mathcal{M}_{decision}$ and compressed models on different mobile devices. The results are shown in Table VI. The results reveal that YOLOv3-tiny exhibits significantly lower latency when compared to deep YOLOv3, which is generally deemed unsuitable for deployment on devices. For instance, the latency of YOLOv3-tiny on Jetson Nano is 87.9% lower than that of YOLOv3. This highlights the substantial potential for accelerating inference using shallow models. It is also evident that $\mathcal{M}_{decision}$ can be executed in real-time on embedded mobile devices such as Jetson Nano, with a latency as low as 23.2 ms, making it suitable for online inference applications.
VII-H Memory and Power Consumptions
We investigate the memory consumption of different models from the following two aspects, i.e., loading model only, and the memory consumption during inference with a batch size of 1. Table VI demonstrates that memory consumption for loading model is significantly lower than that during inference, owing to the presence of hidden parameters during inference. We also examine the impact of different power configurations adopted by Jetson TX2 NX to the performance of Anole. The power consumption and inference speed of Anole and baselines under different power modes are shown in Figure 13, respectively. Anole achieves a 45.1% reduction in power consumption compared with SDM and a inference speed of over 30 FPS with an input power of 20W running 6 cores.
VIII Related Work
VIII-A DNN Prediction on Mobile Devices
To perform DNN inference on mobile devices, new DNNs are specially designed [howard2017mobilenets, zhang2018shufflenet, sanh2019distilbert, jiao2019tinybert] or existing DNNs are compressed to match the computing capability of a mobile device [han2015learning, wang2020pruning]. First, model structure can be optimized to reduce complexity [chen2020addernet, howard2017mobilenets, zhang2018shufflenet]. Second, quantization precision can be reduced to minimize computational cost, e.g., use integers instead of floating-point numbers [alibaba2020mnn, elsen2020fast]. Third, the neural network model can also be accelerated by pruning, i.e., deleting some neurons in the neural network [han2015learning, wang2020pruning, liu2018rethinking]. Scene information is also utilized for model compression on edge/mobile devices [feng2021palleon, xu2022litereconfig, wang2022enabling, jiang2021flexible]. Finally, model distillation can distill the knowledge of large models into small models [sanh2019distilbert, jiao2019tinybert]. Such schemes ensure real-time model inference at the expense of compromised accuracy.
@@ VIII-A DNN Prediction on Mobile Devices
Another direction is to divide DNNs and perform collaborative inference on both edge devices and the cloud [hu2019dynamic, kang2017neurosurgeon, banitalebi2021auto, zheng2022towards, laskaridis2020spinn, guo2021mistify], or to transmit compressed sensory data to the cloud for data recovery and model inference [Liu2019, zhang2021elf]. Neurosurgeon [kang2017neurosurgeon] partitions the computation of each DNN inference task in a layer granularity. CLIO [huang2020clio] addresses the instability of network conditions and optimizes inference under different network states. These approaches need coordination with the cloud for each inference, leading to unpredictable inference delays when the communication link is unstable or disconnected. However, they prove inadequate for cross-scene mobile inference scenarios where even deep models are unable to cope.
VIII-B Cross-scene DNN Prediction
@@
Data-driven machine learning models face challenges in maintaining robust inference performance when dealing with cross-scene inference [zhai2021rise]. One natural approach for scene partitioning is to partition the scene based on prior knowledge or historical samples. First, based on prior knowledge, a similarity graph is constructed to cluster similar domains together [han2014encoding]. However, obtaining such prior knowledge based on domain expertise can be challenging. Second, the original data or their extracted features can be utilized for more automated scene partitioning [lu2022out]. However, these methods may result in the loss of critical information in complex systems [zheng2019metadata]. Cross-scene DNN prediction can also be enhanced given a golden model in the cloud for online sample labelling [khani2023recl, bhardwaj2022ekya]. However, the existence of a perfect or golden model is not always feasible.
VIII-C Mixture of Experts
@@
In recent years, we have witnessed the success of Mixture of Experts (MoE) [jacobs1991adaptive, shazeer2016outrageously], especially in efficient training of large language models (LLM). MoE employs multiple experts for model training, each for one domain. Then, a gate network will be used to determine the correspondence between samples and experts. Though inspired by MoE, Anole differs from MoE in the following 2 aspects. First, experts in MoE are diversified by constraints of losses, but they themselves cannot be related to the scene. In fact, the main purpose of MoE is to expand the number of model parameters, rather than to customize and select scene-specific models. Second, MoE is just a model architecture, and models based on MoE architecture still need to deploy the entire model during deployment. Therefore, MoE-based models often require a significant amount of memory, which is unacceptable for mobile agents like UAVs. In contrast, Anole employs multiple compressed models for online model inference, each designed for one scene. Only a few compressed models are needed to be deployed during online inference. Therefore, Anole is more suitable for mobile devices only with limited resources.
IX Discussion
IX-A How to choose the CMT-DM and CMT-SM algorithm?
The effect of Anole relies on an army of compact DNN models, which can be given in advance or trained by the two algorithms (i.e., Alg.1, CMT-DM and Alg.2, CMT-SM) proposed in Section V-A2. We compare the differences of CMT-DM and CMT-SM, which can guide the choice of the two models training algorithms.
First, in general, CMT-SM is better in terms of inference accuracy than CMT-DM because CMT-SM combines multiple scenes based on both semantics and model training performances, while CMT-SM only considers the similarity of semantics of different scenes. Therefore, with adequate computing resources, we should first choose the CMT-SM algorithm.
However, each combination of scenes requires one round of model training. In fact, the upper bound complexity of the CMT-SM algorithm can be computed by considering the case where no combination of tasks is conducted and the number of scenes never reduces. In this case, the complexity of CMT-SM with $n$ scenes at the $i$th iteration is $O(T_{R}\cdot(n-i+1))$, where $T_{R}$ denotes the time for conducting a training session. Thus, the overall complexity $\mathcal{O}(T_{R}\cdot(\sum_{1}^{n-1}i))\in\mathcal{O}(T_{R}\cdot(n\cdot(n-1)/2))\in\mathcal{O}(T_{R}\cdot n^{2})$. Such a complexity is too large if $T_{R}$ is a large number, e.g., the training time of a typical DNN. In these cases (e.g., the HDR task and the VD task), choosing the CMT-DM algorithm is a practical option.
IX-B Limitation of Anole
Anole has the limitation that, Anole may encounter new scenes which it cannot handle. The performance of Anole relies on a dataset containing all possible scenes and the goal of Anole is to achieve stable model inference across the seen scenes and try our best to generalize to new scenes. The reason why Anole may generalize to unseen new scenes is that the unseen new scenes may be similar to the seen scenes in the view of compressed models. However, we cannot always guarantee this. There may be some new scenes that Anole cannot handle. Fortunately, experiments have shown that such situations are not common. When encountering new scenes that cannot be processed, Anole can make timely reminders and switch to the large model on the cloud to ensure basic inference response.
IX-C Can Mobile Foundation Model (FM) replace Anole?
Mobile Foundation Model [meng2024] inspired by Large Language Model (LLM) such as ChatGPT also proves to be potential to solve the cross-scene online model inference (OMI) problem on mobile devices. Mobile FM also shows the amazing generalization performance on typical mobile tasks. However, Anole can exceed the mobile FM in the following 2 aspects. First, Anole has extremely low memory requirements. Anole only needs to load a cache of compressed models to memory or GPU memory (usually << 2GB), while mobile FM needs to load the whole FM to mobile devices (>> 7.5 GB). As a result, Anole is more suitable for the low-end devices with limited memories. Second, the power consumption of Anole is much less than mobile FM. Despite multiple models are utilized in Anole, each time Anole only use one model for inference. On the contrary, mobile FM needs a whole inference of a large neural network, consuming tens of times more power than Anole. Therefore, Anole is also more suitable for the battery-powered mobile devices such as UAVs. Second, mobile FM cannot work well on all tasks on mobile devices. Without scene-related information, for tasks with less well-annotated data (such as the CGD task), mobile FM cannot solve these task well.
X Conclusion
@@
[Uncaptioned image] Yunzhe Li received his B.S. degree in computer science from Wuhan University in 2021. He is a PhD student in the Department of Computer Science and Engineering, Shanghai Jiao Tong University. His research interests include mobile sensing and edge computing. He has published papers in ICDCS, CIKM, NeurIPS, TMC, TC, etc. He received the Rising Star Award in 2023 from the KubeEdge of Cloud Native Computing Foundation.
[Uncaptioned image] Hongzi Zhu received the PhD degree in computer science from Shanghai Jiao Tong University, in 2009. He was a post-doctoral fellow with the Department of Computer Science and Engineering, Hong Kong University of Science and Technology, and the Department of Electrical and Computer Engineering, University of Waterloo, in 2009 and 2010, respectively. He is now a professor with the Department of Computer Science and Engineering, Shanghai Jiao Tong University. His research interests include mobile sensing, mobile computing, and Internet of Things. He received the Best Paper Award from IEEE Globecom 2016. He is an associate editor for the IEEE Transactions on Vehicular Technology and the IEEE Internet of Things Journal. He is an IEEE Senior Member. For more information, please visit http://lion.sjtu.edu.cn.
[Uncaptioned image] Zhuohong Deng received the B.E. degree in Computer Science and Technology from Beijing Forestry University in China, in 2021. He is currently working toward the Master degree in the Department of Computer Science and Engineering at Shanghai Jiao Tong University. His research interest is artificial intelligence system.
[Uncaptioned image] Yunlong Cheng received the B.E. degree in Computer Science and Technology from Shanghai Jiao Tong University, China, in 2021. Presently, he is pursuing a PhD in Computer Science and Technology at Shanghai Jiao Tong University. His research focuses on data mining, prediction, and scheduling algorithms within the context of cloud services. He has published papers in IPDPS, TCS, DASFAA, etc.
[Uncaptioned image] Zimu Zheng is currently a head research engineer at Huawei Cloud. He received his B.Eng. degree from South China University of Technology and his Ph.D degree from the Hong Kong Polytechnic University. His research interest lies in edge intelligence, benchmarking, multi-task learning, and AIoT. He received the Best Paper Award of ACM e-Energy 2018 and the Best Paper Award of ACM BuildSys 2018. He is an IEEE member with publications in IEEE TPDS, IEEE TPAMI, IEEE IOTJ, IEEE INFOCOM, IEEE ICDCS, IEEE TSUSC, etc. He has also received several awards for outstanding technical contributions in Huawei and serves as a co-chair in the KubeEdge SIG AI of Cloud Native Computing Foundation.
[Uncaptioned image] Liang Zhang received the B.E. degree in Computer Science and Technology from Northeastern University in China, in 2018. She is currently working toward the PhD degree in the Department of Computer Science and Engineering at Shanghai Jiao Tong University. Her research interest includes stream processing and resource scheduling in the cloud or edge computing environment. For more information, please visit https://zl-cs.github.io/.
[Uncaptioned image] Shan Chang received the PhD degree in computersoftware and theory from Xian Jiaotong University,in 2013. From 2009 to 2010, she was a visitingscholar with the Department of Computer Scienceand Engineering, Hong Kong University of Scienceand Technology. She was also a visiting scholar withthe BBCR Research Lab, University of Waterloo,from 2010 to 2011. She is now an associate professor with the Department of Computer Scienceand Technology, Donghua University, Shanghai. Herresearch interests include security and privacy in mobile networks and sensor networks. She is a member of the IEEE.
[Uncaptioned image] Minyi Guo is a Zhiyuan Chair Professor in the Department of Computer Science and Engineering at Shanghai Jiao Tong University, China. He received the PhD degree in computer science from the University of Tsukuba, Japan. He is currently Zhiyuan Chair professor. His present research interests include parallel/distributed computing, compiler optimizations, embedded systems, pervasive computing, big data and cloud computing. He is now on the editorial board for IEEE Transactions on Parallel and Distributed Systems, IEEE Transactions on Cloud Computing and Journal of Parallel and Distributed Computing. He is a fellow of IEEE.